

********************************************************************
UARS WEEKLY STATUS REPORT - 26 Oct 2001
********************************************************************
***********
SOLSTICE
***********
SOLSTICE continues to operate normally, taking solar data on all available
orbital tracking periods and stellar calibration observations whenever
possible. We continue to gather enough data to meet the SOLSTICE science
objectives during this period of reduced tape recorder operations.
***********
SUSIM
***********
SUSIM completed daily mid resolution and MgII solar scans
during periods of predicted UARS telemetry.
***********
ACRIM
***********
Sample ACRIM instrument data was received from PACOR-A for 10/17/01.
Processing software was modified to accomodate the new level0 data
nomenclature. Processing yielded normal results. The ACRIM2 results are
available on the acrim.com web site.
***********
PEM
***********
***** ALL PEM COMMUNICATION SHOULD NOW BE THROUGH THE PEM2K:: NODE *****
The PEM instrument is off at the current time.
PEM would like to thank the project for their work on convincing NASA HQ
to keep UARS alive and operating in campaign mode.
Tests of PEM Level 0 data transfer to PEM2K:: from PACOR-A failed.
Failure was traced to the SwRI firewall blocking file transmission. This has
been corrected and SwRI is ready to again test data transfers from PACOR-A.
*Beginning of power orbit day (PEM activation): array current >5 amps, array
voltage >53.5 volts, and battery current >2.0 amps.
End of power orbit day (PEM deactivation): time in power orbit day >4 minutes,
array current <9 amps, and battery currents <0.2 amps.
Power orbit day variation: ~39-40 min at yaw to ~50-51 min at maximum beta.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
WINDII held its 10th Anniversary Joint Science Team Meeting at
Niagara Falls, Canada, October 22-24. There were 22 attendees, including
not only WINDII colleagues from France, but special guests Ray Roble from
HAO/NCAR, John Emmert from Utah State and Shengpan Zhang from Millstone
Hill. In addition to scientific presentations, plans were made for a
further year of data analysis and operations, consistent with the
proposed Traceability Mission which happily was approved the next day.
***********
MPG
***********
The MPG continues to serve as the point-of-contact
for UARS instrument operations to support re-engineering
of the UARS Mission Operations Concept to develop
capabilities for the Science Traceability Mission.
Daily processing of instrument Daily Activity Plans
continues with no problems.
***********
CDHF
***********
Nothing received.
***********
FOT
***********
REPORTING PERIOD
This report covers 13 October 2001 (Orbit 55156, GMT Day 286) through 19
October 2001 (Orbit 55260, GMT Day 292).
SPACECRAFT OPERATION
The observatory is now in Reverse Flight (Northern Hemisphere viewing) and is
performing nominally under two-battery operations. The instrument operational
changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
HALOE 286/0714 55161 Power OFF
SUSIM 288/0022 55186 Power OFF
SUSIM 292/1316 55254 Power ON (Part 1)
SUSIM 292/1458 55255 Power ON (Complete)
A Forward-to-Reverse Yaw Around maneuver (number 103) was successfully executed
in Orbit 55225 beginning 290/13:55:13 and completing at 290/14:32:23 on 17
October 2001. The SOLSTICE instrument was in day only and all other
instruments were in their off or standby mode (including HRDI and PEM). Due to
the SA position at 269 degrees, the effective night length of the spacecraft
after the maneuver was approximately 56.3 minutes. Post maneuver analysis of
data indicate that during the Yaw Around, the end-of-night load bus voltage was
23.8 volts and the average depths of discharge (DODs) on Batteries 2 and 3 were
approximately 22.1 and 23.4 percent respectively. The next (forward-to-
reverse) YAW is scheduled for 17 October 2001 (DOY 290).
The beta angle ranged between the angles of 19.5 degrees to 7.1 degrees for
this report period. The beta angle is now increasing towards a maximum angle
of 73 degrees on 7 November 2001 (DOY 311).
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery 1
remains OFF the charge relay. Battery 2 and Battery 3 are currently
maintaining greater than 23.8 V end-of-night (EON) load bus voltage. Battery
temperatures are stable with a temperature delta between Battery 2 and Battery
3 of 1.5 to 1.6 degrees C. Additional battery performance data for each
battery on 15 October 2001 and 22 October 2001 are provided in Appendix D.
The clock error ranged from Š15.8 to 13.0 msec during this report period.
Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial excursion of
ALERTS or Out-Of-Limits conditions detected by the FOT for this report. A
summary is listed below and the details of ALERT and Out-Of-Limits occurrences
are listed in Appendix B.
„ OUT-OF-LIMITS = 00
„ ALERTS = 00
The following real-time command sheets were executed for the listed operational
element(s) during this report period (see Appendix A).
„ FOT = 10
„ HALOE = 01
„ SUSIM = 03
„ SOLSTICE = 02
UARS Anomaly Reports initiated during this report period are listed in Appendix
E and summarized below.
„ FOT = 01
Zero (0) Anomaly Reports remained OPEN at the end of this report period (see
Appendix E).
TRANSPONDER FREQUENCIES:
· Transponder A: 01/267 - 2287.496874
· Transponder B: 01/263 - 2287.498962
GROUND SYSTEM OPERATION
One (1) operational support problems occurred this period. Anomaly Reports are
included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from mission
start through UTC Day 01/277 show the observatory data loss to be 303 hours, 18
minutes, and 28.016 seconds. This shows no change in data loss since the last
report period. This is a 0.39 percent data loss, which equals a 99.61 percent
data capture for the mission.
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
The FOT has begun the process of implementing RaytheonÕs Eclipse ground system
and continues to work on the TAC replacement system. These new systems will
provide enhanced capability and make up for the loss of some UARS facilities,
which have been shut down due to budgetary reasons. This will be a huge effort
and must be completed as soon as possible in order to provide a low cost method
for flying UARS. The FOT will need assistance from all entities involved to
insure the success of this operation. The estimated time for completion is
April 2002.
MEETINGS
The UARS FOT attended a UARS Operational Plan Meeting on Wednesday, October 17,
2001.
The UARS FOT attended an Eclipse Training Meeting on Thursday, October 18,
2001.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for 26 November 2001 (DOY 330). This
maneuver will be a Reverse to Forward Flight maneuver (Southern Hemisphere
viewing).
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 19 October 2001
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ ------
55159 286/0529 01-623 NBTR NOOPS D-286
55162 286/0901 01-623 NBTR NOOPS D-286
55168 286/1947 01-623 NBTR NOOPS D-286
55161 286/0714 01-624 HALOE Power OFF
55175 287/0537 01-623 NBTR PBK
55186 288/0022 01-626 SUSIM Power OFF
55186 288/0022 01-627 SOLSTICE DAY-ONLY MODE
55224 290/1304 01-629 MACS YAW MANUEVER START
55225 290/1459 01-629 MACS YAW MANUEVER DONE
55226 290/1638 01-630 MACS TAM FSS
55230 290/2255 01-631 OBC/CLK Clk Adj -20.6 to 0.00 msec/day
55234 291/0444 01-632 OBC/CLK Clk Adj 0.00 to 20.6 msec/day
55246 292/0036 01-633 SOLSTICE Full time ops
55249 292/0400 01-634 OBC Clk Adj fr 20.6 to 0.00 msec/day
55254 292/1316 01-635 SUSIM Power ON (Part 1)
55255 292/1458 01-635 SUSIM Power ON (Complete)
__________________________________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 19 October 2001
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
-------- ----- ----------- ------
-None this report period
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
-------- ----------- ------
-None this report period
_______________________________________________________________________________
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 19 October 2001
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ----- -----------------
55230 290/2302 Clk Adj fr -20.6 to 0.00 msec/day
55234 291/0446 Clk Adj fr 0.00 to 20.6 msec/day
55249 292/0402 Clk Adj fr 20.6 to 0.00 msec/day
___________________________________________________________________________
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 19 October 2001
15 October 2001 - GMT Day 288 Beta = 6.7 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 +728.0/+728.0 -3.68/-4.62 +0.0/ -0.0 11.7 0.000 0.0
2 +00.0/ -95.2 +5.00/+4.19 +28.5/-12.8 25.9 1.040 19.9
3 +22.4/ -22.4 +3.87/+2.92 +27.4/-12.2 25.8 1.020 19.5
22 October 2001 - GMT Day 295 Beta = 21.1 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- ---------- ---------- ----- ------- ------
1 +728.0/+728.0 -3.57/-4.94 +0.0/ -0.0 11.3 0.000 0.0
2 +00.0/ -22.4 +5.15/+3.55 +28.8/-14.5 29.1 1.040 20.9
3 +22.4/ -50.4 +4.19/+2.60 +28.1/-14.3 29.0 1.030 20.5
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- ---- ---------
-no changes V/T 5
_______________________________________________________________________________
APPENDIX E
UARS ANOMALY REPORTS
Week ending 19 October 2001
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
01-060 FOT 55163 286/1123 24071 Drop Out on the Ret. Link
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
01-060 FOT 55163 286/1123 24071 Drop Out on the Ret. Link
ANOMALY INVESTIGATION REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period
SEVNTFW01/42